# Controlled bit-flip of period-doubling and discrete time crystalline states in open systems

### v1

@@ Abstract
In this work, we explore the robustness of a bit-flip operation against thermal and quantum noise for bits represented by the symmetry-broken pairs of the period-doubled (PD) states in a classical parametric oscillator and discrete time crystal (DTC) states in a fully-connected open spin-cavity system, respectively. The bit-flip operation corresponds to switching between the two PD and DTC states induced by a defect in a periodic drive, introduced in a controlled manner by linearly ramping the phase of the modulation of the drive. In the absence of stochastic noise, strong dissipation results in a more robust bit-flip operation in which slight changes to the defect parameters do not significantly lower the success rate of bit-flips. The operation remains robust even in the presence of stochastic noise when the defect duration is sufficiently large. The fluctuations also enhance the success rate of the bit-flip below the critical defect duration needed to induce a switch. By considering parameter regimes in which the DTC states in the spin-cavity system do not directly map to the PD states, we reveal that this robustness is due to the system being quenched by the defect towards a new phase that has enough excitation to suppress the effects of the stochastic noise. This allows for precise control of the bit-flip operations by tuning into the preferred intermediate state that the system will enter during a bit-flip operation. We demonstrate this in a modified protocol based on precise quenches of the driving frequency.
@@ I Introduction
The simplest system that can break discrete time translation symmetry is a single parametric oscillator (PO). When driven resonantly, it enters a period-doubling (PD) state characterized by a subharmonic oscillation relative to the drive accompanied by an exponential growth of the system’s response amplitude [1]. In the case of a nonlinear oscillator, this exponential growth is tapered by the nonlinearity, forcing the response amplitude to relax to a constant value [1]. Similar to systems that spontaneously break discrete symmetries in equilibrium, when a PO enters a PD state, the system randomly chooses one of its two degenerate states, distinguished by a shift of $\pi$ on their oscillation phase [1, 2]. Due to its simplicity, the PO can be emulated in a wide range of setups [3, 4, 5, 6, 2, 7, 8, 9], with the driven pendulum being the simplest example [1]. It has also been extensively used to describe spatio-temporal pattern formation in periodically-driven systems of classical [10, 11, 12, 13] and quantum fluids [14, 15, 16, 17, 18].
@@
Due to the degenerate nature of the PD states, POs have been considered as a good candidate for emulating classical bits [19, 20, 21, 22, 4, 5, 6, 2, 7], motivating the search for robust methods to manipulate these states. It has been shown in Refs. [2, 9, 7] that POs can switch states by varying the natural frequency of the oscillator within a time duration, demonstrating the possibility of performing the simplest classical bit-operation on this system: a bit-flip. For systems with inaccessible natural frequency, however, Ref. [8] has introduced a method for controlled switching between PD states by applying a defect on the driving of the system as shown in Fig. 1(b). The defect protocol consists of a linear ramp of the phase of the drive, $\theta(t)$, from $\theta=0$ to $\theta=2\pi$ within a duration of $T_{\delta}$. Depending on the chosen $T_{\delta}$, the defect protocol can be effectively used as a bit-flip operation for bits encoded in PD states, as shown in Figs. 1(c)-1(g). While Ref. [8] has demonstrated the robustness of this method in noiseless oscillators with weak dissipation, it remains an open question whether it will persist for arbitrary dissipation strength and in the presence of random fluctuations.
@@
Periodically driven systems consisting of multiple constituents can also spontaneously break discrete time translation symmetry [23]. In this case, their collective behavior results in the emergence of a discrete time crystal (DTC). It manifests as a subharmonic oscillation of an order parameter, with an oscillation phase chosen from the multiple degenerate states associated with the broken symmetry [23, 24, 25]. In the case of DTCs with period-doubling oscillations, the system picks an oscillation phase from the two degenerate states owing from the broken $\mathbb{Z}_{2}$ symmetry [24, 25], akin to the PD states of POs. The DTCs have been extensively studied under wide range of platforms, ranging from, but not limited to, networks of classical oscillators [26, 27, 28, 29, 30], spin systems [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 25], Rydberg atoms [41, 42], bosonic systems [43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55], superconductors [56, 57, 58], and particles under oscillating mediums [59, 60, 61].
@@
Figure 1: (a) Sketch of a bit-flip operation. Two degenerate dynamical states, labeled as “0” and “1”, are flipped onto their symmetry-broken pairs by applying a defect protocol within some time duration $T_{\delta}$. (b) (Top panel) Periodic drive with a defect protocol applied during the time window $t=0$ and $t=T_{\delta}=25T_{d}$. (Bottom panel) Phase ramp protocol corresponding to the defect applied in the top panel. (c-h) Sketch and exemplary dynamics of a PO after a defect protocol for (c)-(d) $T_{\delta}<T_{\delta,c}$, (e)-(f) $T_{\delta}=T_{\delta,c}$, and (g)-(h) $T_{\delta}>T_{\delta,c}$. The gray regions correspond to the duration at which the defect protocol is applied. The rest of the parameters are $\omega_{d}=2\Omega$, $A=0.3\Omega$, and $\gamma=0.1\Omega$.
@@
In Ref. [62], it has been shown that there is a correspondence between the degenerate states of the DTCs, which we will refer to as DTC states, in periodically driven fully-connected spin-cavity systems and the PD states of coupled POs. This mapping opens the possibility of manipulating DTC states using the methods for PD states. Indeed, this has been explored for DTC states in other systems, such as closed, integrable bosonic system [45] and classical time crystals [63, 64]. Due to the non-integrability of fully-connected spin-cavity systems, however, such systems generally heat up, leading to a finite lifetime for the DTCs. While this problem can be circumvented by connecting the system to a bath that will absorb the heat generated by the drive [65], it renders the spin-cavity system susceptible to stochastic noise from the environment [66, 67]. This poses the question of whether we can perform bit-flip operations on the DTC states of fully-connected open spin-cavity systems, and how robust the operation would be against quantum fluctuations.
@@
In this work, we address the robustness of bit-flip operations on the PD states of dissipative POs connected to a thermal bath. We then apply these results to demonstrate that robust-bit-flip operations are possible for DTC states of open spin-cavity systems both in the thermodynamic limit and finite-size limit, where quantum fluctuations become dominant. We also consider system parameters leading to DTC states that do not have any direct mapping to PD states. In this case, we demonstrate that the defect protocol can be viewed as a sudden quench towards new dynamical phases. This opens the possibility for more precise control of bit-flip operations on the DTC states.
@@
The paper is structured as follows. In Sec. II, we introduce the systems of interest together with the protocol for the bit-flip. We then present in Sec. III the results for the robustness of the bit-flip operations for arbitrary dissipation without noise and fixed dissipation with arbitrary noise strength. We also demonstrate in this section the applicability of the bit-flip operation on DTC states with no direct mapping to PD states, and give insights on the dynamics during the defect and their implications on the protocol’s robustness against noise. In Sec. IV, we consider other variations of the defect protocol beyond the phase ramp protocol shown in Fig. 1(b) and demonstrate their robustness in performing bit-flips given a set of defect parameters. Finally, we provide a summary and conclusion in Sec. V.
@@ II Systems of interest
Figure 2: Sketch of (a) a dissipative parametric pendulum connected to a thermal bath and (b) a finite-sized open Dicke model. (c)-(f) Zero-temperature bit-flip diagram of the DPP driven at resonance as a function of the defect duration $T_{\delta}$ and amplitude detuning $\delta A=A-A_{r}$ for increasing dissipation strength, (c) $\gamma=0.005\Omega$, (d) $\gamma=0.01\Omega$, (e) $\gamma=0.025\Omega$, and (f) $\gamma=0.1\Omega$. (g)-(j) Bit-flip diagram of the resonantly-driven ODM in the thermodynamic limit for (g) $\kappa=0.01\omega$, (h) $\kappa=0.02\omega$, (i) $\kappa=0.06\omega$, (j) $\kappa=0.1\omega$. The remaining parameters for the ODM are $\lambda_{0}=0.9\lambda_{c}$, and $\omega_{0}=\omega$.
As a testbed for exploring the robustness of bit-flips on PD states for noisy POs, we consider a parametrically driven pendulum of mass $m$ and length $L$ coupled to a thermal bath, depicted in Fig. 2(a). The system is described by the Langevin equation [68]
@@
In the zero-temperature limit, $\tilde{T}\rightarrow 0$, the stochastic noise is absent, and thus Eq. (1) reduces to the equations of motion for a dissipative parametric pendulum (DPP). The DPP can be driven resonantly by setting $\omega_{d}$ and $A$ to [1, 20] (see Appendix B.1 for more details),
$\omega_{r,\mathrm{DPP}}=2\Omega,\quad A_{r,\mathrm{DPP}}=2\gamma/\Omega,$ (5)
where $\omega_{r,\mathrm{DPP}}$ is the primary resonant frequency of the DPP and $A_{r,\mathrm{DPP}}$ is the resonant amplitude corresponding to the minimum $A$ needed to enter a PD state at $\omega_{r,\mathrm{DPP}}$. In this case, the system exhibits a period-doubled oscillation with respect to the driving period, $T_{d}$, with an oscillation phase spontaneously chosen from the two degenerate PD states [1, 2]. This phase shift, which we refer to as the absolute-time phase $\varphi$, can be extracted from an order parameter, $O(t)$, capturing the system’s period-doubling response using the equation [63]
@@
To test the bit-flip operations for the DTC states, we consider the open Dicke model (ODM) described by the Heisenberg-Langevin equation [69]
$\frac{\partial}{\partial t}\hat{a}=\frac{i}{\hbar}\left[\hat{H},\hat{a}\right]-\kappa\hat{a}+\xi(t),$ (7a)
$\frac{\partial}{\partial t}\hat{S}^{x,y,z}=\frac{i}{\hbar}\left[\hat{H},\hat{S}^{x,y,z}\right],$ (7b)
@@
[…]
$\xi(t)\right>=0,\quad\left<\xi(t)\xi(t^{\prime})\right>=\kappa\delta\left(t-t^{\prime}\right).$ (9)
Throughout this paper, we set $\omega=\omega_{0}$.
We can establish the noiseless limit of the ODM by either setting $\kappa=0$ or considering the thermodynamic limit, $N\rightarrow\infty$. In the latter case, we introduce the rescaled quantities $\hat{\alpha}=\hat{a}/\sqrt{N}$, $\hat{s}^{x,y,z}=\hat{S}^{x,y,z}/N$. Substituting these to the equations of motion derived from Eq. (7) results in the renormalization of the stochastic noise from $\xi(t)$ into $\tilde{\xi}(t)=\xi(t)/\sqrt{N}$ (see details in Appendix A). By setting $N\rightarrow\infty$, the contributions of the quantum fluctuations become negligible in the thermodynamic limit, thereby the dynamics is exactly solvable by mean-field analysis [72], in which we set $\hat{a}\rightarrow a\in\mathbb{C}$ and $\hat{S}^{x,y,z}\rightarrow S^{x,y,z}\in\mathbb{R}$.
In the thermodynamic and static limit, $\lambda(t)\rightarrow\lambda_{0}$, the ODM has two equilibrium phases: the normal phase (NP) and the superradiant phase (SP) [70, 71]. For coupling strengths less than a critical value, $\lambda_{0}<\lambda_{c}$, the system prefers the NP characterized by a fully polarized collective spin $S^{z}$ at the $-z$ direction, and zero photon number. On the other hand, for $\lambda_{0}>\lambda_{c}$, the system enters the SP, which manifests in the cavity having a nonzero photon number and a nonzero $S^{x}$ component, the direction of which is chosen from the two degenerate steady states of the system [71]. The critical point separating the NP and the SP is
$\lambda_{c}=\frac{1}{2}\sqrt{\kappa^{2}+\omega^{2}}.$ (10)
@@
We investigate the bit-flip operation on the PD states of the DPP and the DTC states of the ODM by using a defect protocol based on a linear ramp of the phase of the drive. In particular, given an arbitrary periodic drive with a time-dependent phase $\theta(t)$
@@
where $T_{\delta}$ is the defect duration, and $t_{i}$ and $t_{f}$ are the initial and final time, respectively. Throughout this work, we consider the $x$-quadrature of the DPP, $x=\sin(u)$, as its main order parameter, while we use the $x$-component of the collective spin operator, $S^{x}$, for the ODM.
In the limit of $T_{\delta}\gg T_{d}$, the defect protocol in Eq. (14) will always induce a bit flip on the PD and DTC states since the absolute-time phase of both states adiabatically follow $\theta(t)$ when the ramp is sufficiently slow [8]. In particular, during the defect protocol, the driving frequency effectively becomes $\omega_{d}^{\prime}=\omega_{d}+2\pi/T_{\delta}$ for any $T_{\delta}$. Since the order parameter of both the DPP and the ODM oscillates at half the driving frequency, we can infer that during the defect protocol for large $T_{\delta}$,
@@
where $\varphi_{0}$ is the absolute-time phase before the defect. Thus at $t\geq T_{\delta}$, both systems acquire a phase shift of $\pi$, leading to a perfect bit-flip success rate for $T_{\delta}\gg T_{d}$. In the following section, we are interested in the response of the system and the robustness of the bit-flip operations when $T_{\delta}$ is close to $T_{d}$.
III PD and DTC state switching

### v2

# Controlled bit flip of period-doubling and discrete time crystalline states in open systems

@@ Abstract
In this work, we explore the robustness of a bit-flip operation against thermal and quantum noise for bits represented by the symmetry-broken pairs of the period-doubled states in a classical parametric oscillator and discrete time crystal (DTC) states in an open fully-connected spin-cavity system, respectively. The bit-flip operation corresponds to switching between the two period-doubled and DTC states induced by a defect in a periodic drive. The defect is introduced in a controlled manner by linearly ramping the phase of the modulation of the drive. In the absence of stochastic noise, strong dissipation results in a more robust bit-flip operation in which slight changes to the defect parameters do not significantly lower the success rate of bit-flips. The operation remains robust even in the presence of stochastic noise when the defect duration is sufficiently large. By considering parameter regimes in which the DTC states in the spin-cavity system do not directly map to the period-doubled states, we reveal that this robustness stems from the system being quenched by the defect towards a new phase that has enough excitation to suppress the effects of the stochastic noise. This allows for precise control of the bit-flip operations by tuning into the preferred intermediate state that the system will enter during a bit-flip operation. We demonstrate this in a modified protocol based on precise quenches of the driving frequency.
@@ I Introduction
The simplest system that can break discrete time translation symmetry is a single parametric oscillator (PO). When driven resonantly, it enters a period-doubled state characterized by a subharmonic oscillation relative to the drive accompanied by an exponential growth of the system’s response amplitude [1]. In the case of a nonlinear oscillator, this exponential growth is tapered by the nonlinearity, forcing the response amplitude to relax to a constant value [1]. Similar to systems that spontaneously break discrete symmetries in equilibrium, when a PO enters a period-doubled state, the system randomly chooses one of its two degenerate states, distinguished by a shift of $\pi$ on their oscillation phase [1, 2]. Due to its simplicity, the PO can be emulated in a wide range of setups [3, 4, 5, 6, 2, 7, 8, 9], with the driven pendulum being the simplest example [1].
@@
Due to the degenerate nature of the period-doubled states, POs have been considered as a good candidate for emulating classical bits [10, 11, 12, 13, 4, 5, 6, 2, 7], motivating the search for robust methods to manipulate these states. It has been shown in Refs. [2, 9, 7] that POs can switch states by varying the natural frequency of the oscillator within a time duration, demonstrating the possibility of performing the simplest classical bit-operation on this system: a bit-flip. For systems with inaccessible natural frequency, however, Ref. [8] has introduced a method for controlled switching between period-doubled states by applying a defect on the system’s drive as shown in Fig. 1(b). The defect protocol consists of linearly ramping the phase of the drive, $\theta(t)$, from $\theta=0$ to $2\pi$ within a duration of $T_{\delta}$. Depending on the chosen $T_{\delta}$, the defect protocol can be effectively used as a bit-flip operation for bits encoded in the period-doubled states, as shown in Figs. 1(c)-1(g). While Ref. [8] has demonstrated the robustness of this method in noiseless oscillators with weak dissipation, it remains an open question whether it will persist for arbitrary dissipation strength and in the presence of random fluctuations.
@@
Periodically driven many-body systems can also spontaneously break discrete time translation symmetry [14]. In this case, their collective behavior results in the emergence of a discrete time crystal (DTC). It manifests as a subharmonic oscillation of an order parameter, with an oscillation phase chosen from the multiple degenerate states associated with the broken symmetry [14, 15, 16]. In the case of DTCs with period-doubling oscillations, the system picks an oscillation phase from the two degenerate states owing from the broken $\mathbb{Z}_{2}$ symmetry [15, 16], akin to the period-doubled states of POs. The DTCs have been extensively studied under wide range of platforms, ranging from, but not limited to, networks of classical oscillators [17, 18, 19, 20, 21], spin systems [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 16], Rydberg atoms [32, 33], bosonic systems [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46], superconductors [47, 48, 49], and particles under oscillating mediums [50, 51, 52].
@@
Figure 1: (a) Sketch of a bit-flip operation. Two degenerate dynamical states, labeled as “0” and “1”, are flipped onto their symmetry-broken pairs by applying a defect protocol within some time duration $T_{\delta}$. (b) (Top panel) Periodic drive with a defect protocol applied during the time window $t=0$ and $t=T_{\delta}=25T_{d}$. (Bottom panel) Phase ramp protocol corresponding to the defect applied in the top panel. (c)-(h) Sketch and exemplary dynamics of a PO after a defect protocol for (c), (d) $T_{\delta}<T_{\delta,c}$, (e), (f) $T_{\delta}=T_{\delta,c}$, and (g), (h) $T_{\delta}>T_{\delta,c}$. The gray regions correspond to the duration at which the defect protocol is applied. The rest of the parameters are $\omega_{d}=2\Omega$, $A=0.3\Omega$, and $\gamma=0.1\Omega$.
@@
In Ref. [53], it has been shown that there is a correspondence between the degenerate states of the DTCs, which we will refer to as DTC states, in periodically driven fully-connected spin-cavity systems and the period-doubled states of coupled POs. This mapping opens the possibility of manipulating DTC states using the methods for period-doubled states. Indeed, this has been explored for DTC states in other systems, such as closed, integrable bosonic system [36] and classical time crystals [54, 55]. Due to the non-integrability of fully-connected spin-cavity systems, however, such systems generally heat up, leading to a finite lifetime for the DTCs. While this problem can be circumvented by connecting the system to a bath that will serve as a heat sink [56], it renders the spin-cavity system susceptible to stochastic noise from the environment [57, 58, 59]. This poses the question of whether we can perform bit-flip operations on the DTC states of fully-connected open spin-cavity systems, and how robust the operation would be against quantum fluctuations.
@@
In this work, we address the robustness of bit-flip operations on the period-doubled states of POs connected to a thermal bath. We then apply these results to demonstrate that robust-bit-flip operations are possible for DTC states of open spin-cavity systems both in the thermodynamic limit and finite-size limit, where quantum fluctuations become dominant. We also consider system parameters leading to DTC states that do not have any direct mapping to period-doubled states. In this case, we demonstrate that the defect protocol can be viewed as a sudden quench towards new dynamical phases. This opens the possibility for more precise control of bit-flip operations on the DTC states.
@@
The paper is structured as follows. In Sec. II, we introduce the systems of interest together with the protocol for the bit-flip. We then present in Sec. III the results for the robustness of the bit-flip operations for arbitrary dissipation without noise, and fixed dissipation with arbitrary noise strength. We also demonstrate in this section the applicability of the bit-flip operation on DTC states with no direct mapping to period-doubled states, and give insights on the dynamics during the defect and their implications on the protocol’s robustness against noise. In Sec. IV, we consider other variations of the defect protocol beyond the phase ramp protocol shown in Fig. 1(b) and demonstrate their robustness in performing bit-flips given a set of defect parameters. Finally, we provide a summary and conclusion in Sec. V.
@@ II Systems of interest
Figure 2: Sketch of (a) a parametric oscillator connected to a thermal bath and (b) a finite-sized open Dicke model. (c)–(f) Zero-temperature bit-flip diagram of the PO driven at resonance as a function of the defect duration $T_{\delta}$ and amplitude detuning $\delta A=A-A_{r}$ for increasing dissipation strength, (c) $\gamma=0.005\Omega$, (d) $\gamma=0.01\Omega$, (e) $\gamma=0.025\Omega$, and (f) $\gamma=0.1\Omega$. (g)-(j) Bit-flip diagram of the resonantly driven ODM in the thermodynamic limit for (g) $\kappa=0.01\omega$, (h) $\kappa=0.02\omega$, (i) $\kappa=0.06\omega$, (j) $\kappa=0.1\omega$. The remaining parameters for the ODM are $\lambda_{0}=0.9\lambda_{c}$, and $\omega_{0}=\omega$.
As a test bed for exploring the robustness of bit-flips on period-doubled states, we consider a PO of mass $m$ and length $L$ coupled to a thermal bath, depicted in Fig. 2(a). The system is described by the Langevin equation [60]
@@
In the zero-temperature limit, $\tilde{T}\rightarrow 0$, the PO can be driven resonantly by setting $\omega_{d}$ and $A$ to [1, 11] (see Appendix B.1 for more details),
$\omega_{r,\mathrm{PO}}=2\Omega,\quad A_{r,\mathrm{PO}}=2\gamma/\Omega,$ (5)
where $\omega_{r,\mathrm{PO}}$ is the primary resonant frequency of the PO and $A_{r,\mathrm{PO}}$ is the minimum $A$ needed to enter a period-doubled state at $\omega_{r,\mathrm{PO}}$. In this case, the system exhibits a period-doubled oscillation with respect to the driving period, $T_{d}$, with an oscillation phase spontaneously chosen from the two degenerate period-doubled states [1, 2]. This phase shift, which we refer to as the absolute-time phase $\varphi$, can be extracted from an order parameter, $O(t)$, capturing the system’s period-doubling response using the equation [54]
@@
To test the bit-flip operations for the DTC states, we consider the open Dicke model (ODM) described by the Lindblad master equation [61]
$\partial_{t}\hat{\rho}=-i\left[\hat{H}/\hbar,\hat{\rho}\right]+\kappa\left(2\hat{a}\hat{\rho}\hat{a}^{\dagger}-\left\{\hat{a}^{\dagger}\hat{a},\hat{\rho}\right\}\right),$ (7)
@@
The ODM, depicted in Fig. 2(b), describes the interaction of $N$ two-level systems coupled to a single cavity mode that is connected to a Markovian bath [62, 61]. The two-level systems are represented by the collective spin operators $\hat{S}^{x,y,z}=\sum_{i=1}^{N}\hat{\sigma}_{i}^{x,y,z}/2$, with $\sigma_{i}^{x,y,z}$ representing the $i$-th two-level system. Meanwhile, the cavity mode is represented by the bosonic creation (annihilation) operator $\hat{a}^{\dagger}$ ($\hat{a}$). The transition frequencies of the cavity and spins are $\omega$ and $\omega_{0}$, respectively. The time-dependent spin-cavity coupling is $\lambda(t)=\lambda_{0}f(t)$. Throughout this work, we consider $\omega=\omega_{0}$.
In the thermodynamic limit, $N\rightarrow\infty$, and static limit, $\lambda(t)\rightarrow\lambda_{0}$, the ODM has two equilibrium phases: the normal phase (NP) and the superradiant phase (SP) [62, 61]. For coupling strengths less than a critical value, $\lambda_{0}<\lambda_{c}$, the system prefers the NP characterized by a fully polarized collective spin $S^{z}\equiv\langle\hat{S}_{z}\rangle$ at the $-z$ direction, and zero photon number. On the other hand, for $\lambda_{0}>\lambda_{c}$, the system enters the SP, which manifests as having a nonzero photon number and a nonzero $S^{x}\equiv\langle\hat{S}^{x}\rangle$ component, the direction of which is chosen from the two degenerate steady states of the system [61]. The critical point separating the NP and the SP is $\lambda_{c}=\frac{1}{2}\sqrt{\kappa^{2}+\omega^{2}}.$
@@
We investigate the bit-flip operation on the period-doubled and DTC states using a defect protocol based on a linear ramp of the phase of the drive. In particular, given a periodic drive with a time-dependent phase $\theta(t)$
@@
where $T_{\delta}$ is the defect duration, and $t_{i}$ and $t_{f}$ are the initial and final time, respectively. Throughout this work, we consider the $x$-quadrature of the PO, $x=\sin(u)$, as its main order parameter, while we use $S^{x}$ for the ODM.
In the limit of $T_{\delta}\gg T_{d}$, the defect protocol in Eq. (12) will always induce a bit flip on the period-doubled and DTC states since the absolute-time phase of both states adiabatically follow $\theta(t)$ when the ramp is sufficiently slow [8]. In particular, during the defect protocol, the driving frequency effectively becomes $\omega_{d}^{\prime}=\omega_{d}+2\pi/T_{\delta}$ for any $T_{\delta}$. Since the order parameter of both the PO and the ODM oscillates at half the driving frequency, we can infer that during the defect protocol for large $T_{\delta}$,
@@
where $\varphi_{0}$ is the absolute-time phase before the defect. Thus at $t\geq T_{\delta}$, both systems acquire a phase shift of $\pi$, leading to a perfect bit-flip success rate for $T_{\delta}\gg T_{d}$. In the following section, we are interested in the response of the two systems and the robustness of the bit-flip operations when $T_{\delta}$ is close to $T_{d}$.
III Period-doubled and DTC state switching